Use simple interest to find the ending balance, given the following values for principal, rate, and time. 9) $34,000 at 11.8% for 5 years
Answer #1

9)

P=34000,r=11.8\%,t=5 years


total interest ,i=Prt
=>i=34000*(11.8/100)*5
=>i=\$ 20060

ending balance =P+i
=>ending balance =34000+20060
=>ending balance =\$54060

PLEASE RATE IF HELPFUL.PLEASE COMMENT IF YOU HAVE ANY DOUBT